# Quantifying Resolution Limits in Pedestal Profile Measurements with Gaussian Process Regression

### v1

@@ 6 Summary and Outlook
In summary, this work has developed two tools, a low-pass filter analogy and $N_{eff}$, to quantify resolution limits for the inference of differentiable profiles using GPR. These tools can be used to replace trial-and-error in the fitting process with quantitative metrics that can be used to engineer the tradeoff between over-fitting and over-regularization. We summarize the main practical takeaways from the analyses:
  • For finitely sampled data, GPR does not infer the ‘true’ ensemble averaged profile from the data, but rather a low-pass filtered version of this profile. Examples in section 5 show how over-regularization can reduce gradients by $\sim 10\%$ and conceal ‘flat spots’ associated with low-order rational surfaces. Given the tendency for profiles to lie near marginal stability boundaries, this low-pass filtering could lead to erroneous conclusions about predicted transport or the stability of different modes.
@@
  • A procedure for determining estimates or bounds on hyperparameters is given in section 3.3. For inferring profiles with rapid changes in the derivative, the squared exponential kernel may be vulnerable to ‘ringing’ artifacts associated with the Gibbs phenomenon; the Matérn kernel with $\nu=2.5$ can be considered as an alternative. To avoid over-regularization, $\ell$ and $S$ should be chosen so that $\xi_{*}w_{ped}\lesssim 1$, where $w_{ped}$ is the full width at half maximum of the inferred pedestal profile.
@@
  • The $N_{eff}$ parameter can be used to quantify the credibility of a given GPR inference in terms of the number of independent measurements contributing to the inference a given point. It does not rely on any simplifying assumptions and can be evaluated for any choice of kernel. A higher $N_{eff}$ indicates a higher robustness to unmodeled errors in the inference reducing the risk of overfitting, although this usually also implies stronger regularization of the fitted profiles.
@@
Detailed demonstrations of how this advice can be applied in practice for inferring both time-independent and time-dependent pedestal profiles are shown in section 5. In particular, we demonstrate how analyzing the low-pass filtering associated with GPR gives a quantitative understanding of the smoothing applied by the inference, leading to better estimates of the profile gradients. We also demonstrate the power of GPR to credibly infer time-dependent evolution of pedestal properties over an inter-ELM cycle while simultaneously inferring subtle features such as ‘flat spots’ associated with low-order rational surfaces.
@@
Finally, we remark on current limitations and possible extensions of this work. The analysis here was primarily done assuming spatially localized measurements with a known, independent Gaussian noise. Both of these assumptions can be relaxed somewhat. As mentioned in the introduction, several methods for GPR tomography and integrating data from multiple heterogeneous diagnostic sources exist. Furthermore, correlated measurement error can be directly encoded into the covariance matrix $\boldsymbol{\Sigma}_{\varepsilon}$, and there exist methods to infer noise levels from measurements [49]. We also remark that we have primarily explored GPR without the addition of constraints, such as monotonicity or non-negativity of the inferred profiles. Different methods have been proposed for imposing these constraints, such as monotonicity promoting priors [23] or enforcing linear inequality constraints [50]. As long as an explicit linear or affine relationship between the measurements and the inferred profile can be computed, the equivalent kernel analysis can likely be extended to these cases to derive similar quantities like the cutoff frequency and $N_{eff}$.
@@ Appendix A Derivation of Low-Pass Filter
̊$\mathring{\mathcal{F}}[g_{j}](\xi\Delta x)=\sum_{j=-\infty}^{\infty}g_{j}e^{-\mathfrak{i}2\pi\xi\Delta xj}$
Note the continuous and discrete Fourier transforms are related [51] by
̊$\mathring{\mathcal{F}}[f(x_{i})](\xi\Delta x)=\frac{1}{\Delta x}\sum_{j=-\infty}^{\infty}\mathcal{F}[f](\xi+\tfrac{j}{\Delta x})$
@@
̊$\mathring{\mathcal{F}}[f(x_{i})](\xi\Delta x)\approx\frac{1}{\Delta x}\mathcal{F}[f](\xi)$
@@ Appendix B Derivation of $N_{eff}$
To compute the Fisher information of $\mathbf{Y}$ with respect to $\bar{f}(x)$, which is a linear combination of the parameters $f(\mathbf{x})$, we need to marginalize out the other $N-1$ degrees of freedom from the Fisher information matrix. This can be accomplished by finding a transformation of our parameters $f(\mathbf{x})$ such that the information matrix with respect to the new parameter coordinates will be diagonal [52].
@@ Appendix C Derivation of $N_{eff}$ Bound
To derive inequality (16), we use the Cauchy-Schwarz inequality by treating the variance in the numerator as the inner product between two vectors $(\sigma_{1}^{2}\beta_{1}(x)^{2},...,\sigma_{N}^{2}\beta_{N}(x)^{2})$ and $(\delta_{1}^{2},...,\delta_{N}^{2})$,
$\operatorname{Var}[\tilde{f}(x)-\hat{f}(x)]^{2}=\left(\sum_{i=1}^{N}\sigma_{i}^{2}\beta_{i}(x)^{2}\delta_{i}^{2}\right)^{2}\leq\left(\sum_{i=1}^{N}\sigma_{i}^{4}\beta_{i}(x)^{4}\right)\left(\sum_{i=1}^{N}\delta_{i}^{4}\right)$
Dividing both sides by $\bar{\sigma}_{\varepsilon}^{2}(x)$ results in the desired expression.

### v2

@@ 6 Summary and Outlook
In summary, this work has developed two tools, a low-pass filter analogy and $N_{eff}$, to quantify resolution limits for the inference of differentiable profiles using GPR. These tools can be used to replace trial-and-error in the fitting or hyperprior development process with quantitative metrics that can be used to engineer the tradeoff between over-fitting and over-regularization. We summarize the main practical takeaways from the analyses:
  • For finitely sampled data, GPR does not infer the ‘true’ ensemble averaged profile from the data, but rather a low-pass filtered version of this profile. Examples in section 5 show how over-regularization can reduce gradients on the order of $\sim 10\%$ compared to properly regularized profiles. Given the tendency for profiles to lie near marginal stability boundaries, this low-pass filtering could impact the predicted transport or the stability of different modes.
@@
  • A procedure for determining estimates or bounds on hyperparameters is given in section 3.3. For inferring profiles with rapid changes in the derivative, the squared exponential kernel may be vulnerable to ‘ringing’ artifacts associated with the Gibbs phenomenon; the Matérn kernel with $\nu=2.5$ can be considered as an alternative. To avoid over-regularization, $\ell$ and $S$ should be chosen so that $\xi_{*}w_{ped}\gtrsim 1$, where $w_{ped}$ is the full width at half maximum of the inferred pedestal profile.
@@
  • The $N_{eff}$ parameter quantifies the number of independent measurements contributing to the inference a given point, which allows for the credibility of GPR to be assessed in regards to over-fitting. It does not rely on any simplifying assumptions and can be evaluated for any choice of kernel. A higher $N_{eff}$ indicates a higher robustness to unmodeled errors in the inference reducing the risk of over-fitting, although this usually also implies stronger regularization of the fitted profiles.
  • By computing low-pass filter cutoffs and $N_{eff}$ across ranges of hyperparameter space, it is possible to use GPR to compute point estimate inferences with precisely controlled levels of regularization. Although the error bars provided by these inferences will only capture epistemic (within-model) uncertainty and not aleatoric (between-model) uncertainty, the inferences do represent unbiased estimates of approximately low-pass filtered versions of the ‘true’ profiles. By cross-referencing against physical expectations, e.g. alignment with magnetic safety factor profiles, it is possible to use these metrics to build credibility in the inference of features with increasingly fine spatial and temporal scales.
@@
Detailed demonstrations of how this advice can be applied in practice are given in a case study inferring both time-independent and time-dependent pedestal profiles are shown in section 5. In particular, we demonstrated how analyzing the low-pass filtering associated with GPR gives a quantitative understanding of the smoothing applied by the inference, leading to a better understanding of credible interpretations of pedestal measurements into quantities such as the peak gradient and pedestal width. We also demonstrated the power of GPR to credibly infer time-dependent evolution of pedestal properties over an inter-ELM cycle while simultaneously inferring subtle features such as ‘flat spots’ associated with low-order rational surfaces.
@@
Finally, we remark on current limitations and possible extensions of this work. The analysis here was primarily done assuming spatially localized measurements with a known, independent Gaussian noise. Both of these assumptions can be relaxed somewhat. As mentioned in the introduction, several methods for GPR tomography and integrating data from multiple heterogeneous diagnostic sources exist. Furthermore, correlated measurement error can be directly encoded into the covariance matrix $\boldsymbol{\Sigma}_{\varepsilon}$, and there exist methods to infer noise levels from measurements [mathews_quantifying_2021]. We also remark that we have primarily explored GPR without the addition of constraints, such as monotonicity or non-negativity of the inferred profiles. Different methods have been proposed for imposing these constraints, such as monotonicity promoting priors [michoski_gaussian_2024] or enforcing linear inequality constraints [da_veiga_gaussian_2012]. As long as an explicit linear or affine relationship between the measurements and the inferred profile can be computed, the equivalent kernel analysis can likely be extended to these cases to derive similar quantities like the cutoff frequency and $N_{eff}$.
@@
Data Availability Statement
The code associated with this work can be found at https://github.com/Maplenormandy/gpr-resolution
@@ Appendix A Derivation of Low-Pass Filter
$\mathring{\mathcal{F}}[g_{j}](\xi\Delta x)=\sum_{j=-\infty}^{\infty}g_{j}e^{-\mathfrak{i}2\pi\xi\Delta xj}$
Note the continuous and discrete Fourier transforms are related [proakis_digital_1996] by
$\mathring{\mathcal{F}}[f(x_{i})](\xi\Delta x)=\frac{1}{\Delta x}\sum_{j=-\infty}^{\infty}\mathcal{F}[f](\xi+\tfrac{j}{\Delta x})$
@@
$\mathring{\mathcal{F}}[f(x_{i})](\xi\Delta x)\approx\frac{1}{\Delta x}\mathcal{F}[f](\xi)$
@@ Appendix B Derivation of $N_{eff}$
To compute the Fisher information of $\mathbf{Y}$ with respect to $\bar{f}(x)$, which is a linear combination of the parameters $f(\mathbf{x})$, we need to marginalize out the other N1N-1 degrees of freedom from the Fisher information matrix. This can be accomplished by finding a transformation of our parameters $f(\mathbf{x})$ such that the information matrix with respect to the new parameter coordinates will be diagonal [barndorff-nielsen_inference_2017].
@@ Appendix C Derivation of $N_{eff}$ Bound
For convenience, we reiterate the formula
$\tilde{f}(x)-\hat{f}(x)=\sum_{j=1}^{M}\left[\delta_{j}\sum_{i\in P_{j}}\beta_{i}(x)\sigma_{i}\right]$ (27)
and notice that the derivative satisfies a similar equation
$\tilde{f}^{\prime}(x)-\hat{f}^{\prime}(x)=\sum_{j=1}^{M}\left[\delta_{j}\sum_{i\in P_{j}}\beta_{i}^{\prime}(x)\sigma_{i}\right]$
Inequality (19) is derived by the following steps,
$\operatorname{Var}[\tilde{f}(x)-\hat{f}(x)|f]^{2}$ $=\left(\sum_{j=1}^{M}\operatorname{Var}[\delta_{j}]\left(\sum_{i\in P_{j}}\sigma_{i}\beta_{i}(x)\right)^{2}\right)^{2}$
$\leq\left(\sum_{j=1}^{M}|P_{j}|\operatorname{Var}[\delta_{j}]\sum_{i\in P_{j}}\sigma_{i}^{2}\beta_{i}(x)^{2}\right)^{2}$
$\leq\left(\sum_{j=1}^{M}|P_{j}|^{2}\operatorname{Var}[\delta_{j}]^{2}\right)\left(\sum_{j=1}^{M}\left(\sum_{i\in P_{j}}\sigma_{i}^{2}\beta_{i}(x)^{2}\right)^{2}\right)$
$=\overline{D^{2}}\frac{\sigma_{\varepsilon}^{2}(x)^{2}}{N_{\mathbf{P}}[\bar{f}(x)]}$
where dividing both sides by $\sigma_{\varepsilon}^{2}(x)^{2}$ gives the desired result.
To explain this sequence of steps, the first line uses the fact that the $\delta_{j}$ are all independent, and hence the variance of the sum (27) is the sum of the variances, and also uses $\operatorname{Var}[\alpha X]=\alpha^{2}\operatorname{Var}[X]$ for constant $\alpha$. The second line uses the relationship between the arithmetic mean and the root mean square,
$\left(\sum_{i\in P_{j}}\sigma_{i}\beta_{i}(x)\right)^{2}\leq|P_{j}|\sum_{i\in P_{j}}\sigma_{i}^{2}\beta_{i}(x)^{2}$
where equality holds if all of the $\sigma_{i}\beta_{i}(x)$ are equal for $i\in P_{j}$.
The third line comes from an application of the Cauchy-Schwarz inequality $(\mathbf{a}\cdot\mathbf{b})^{2}\leq\left\lVert a\right\rVert^{2}\left\lVert b\right\rVert^{2}$ on two vectors $\mathbf{a},\mathbf{b}$ which have components $a_{j}=|P_{j}|\operatorname{Var}[\delta_{j}]$ and $b_{j}=\sum_{i\in P_{j}}\sigma_{i}^{2}\beta_{i}(x)^{2}$. Notice that $\mathbf{a}\cdot\mathbf{b}=\left\lVert a\right\rVert\left\lVert b\right\rVert\cos(\theta)$ where $\theta$ is the angle between $\mathbf{a}$ and $\mathbf{b}$, so equality holds when $\mathbf{a}\parallel\mathbf{b}$.
The fourth and final line comes from
$\frac{1}{N_{\mathbf{P}}[\bar{f}(x)]}=\frac{\sum_{j=1}^{M}\left(\sum_{i\in P_{j}}\sigma_{i}^{2}\beta_{i}(x)^{2}\right)^{2}}{\left(\sum_{k=1}^{N}\sigma_{k}^{2}\beta_{k}(x)^{2}\right)^{2}}$
Finally, notice that if $\beta_{i}(x)$ is replaced by $\beta_{i}^{\prime}(x)$ in each line of the derivation, an exactly analogous result is derived for $\tilde{f}^{\prime}(x)-\hat{f}^{\prime}(x)$.
From the discussion earlier, we can show that the worst-case scenario predicted by the bound can happen in the following scenario: Suppose for each bin $P_{j}$, each measurement $i\in P_{j}$ has the same error bar $\sigma_{i}$ and measurement location $x_{i}$, which leads to all of the $\sigma_{i}\beta_{i}(x)$ being equal in the bin and hence equality on the second line. Then, the components $b_{j}=|P_{j}|\sigma_{i}^{2}\beta_{i}(x)^{2}$, so if $\operatorname{Var}[\delta_{j}]\propto\sigma_{i}^{2}\beta_{i}(x)^{2}$ for one of the $i\in P_{j}$, then equality holds on the third line. This achieves the worst-case deviation predicted by the bound. Intuitively, this represents the case where the variance of the unmodeled error $\delta_{j}$ is proportional to the squared sensitivity of the inference to the measurements in the bin $P_{j}$.